A helium-filled balloon at sea level has a volume of 1.5 L at 0.979 atm and 35°C. If it is released and rises to an elevation at which the pressure is 0.900 atm and the temperature is 25°C, what will be the new volume of the balloon?
1
Expert's answer
2017-02-14T04:37:18-0500
Given
p0 = 0.979 atm
v0 = 1.5 L
t0 = 35 °C

p1 = 0.9 atm
t1 = 25 °C
What asling for v1
Solution
According to Ideal Gas Low
p0 * V0 / t0 = p1 * V1 / t1
Where from
v1 = (p0 * V0 / t0) *t1 /p1 = 1.17 L